Zebra has a total fixed cost of $240,000 and a variable cost of $40/unit. They are planning to hire a CFO as their business has grown. Determine the salary the company can pay the CFO if they have determined that they can sell 30,000 units at a price of $65 and desire to earn a profit of $250,000.
Answer #1

Let the salary be x

Total Contribution - Fixed cost - salary = Net incone

[30,000*(65-40)] - 240,000 - salary = 250,000

750,000 - 240,000 - salary = 250,000

Salary = 750,000-240,000-250,000

= 260,000

Thd company can pay 260,000 as salary